OpenStudy (anonymous):

Circle S has an equation of (x – 16)2 + (y – 9)2 = 4. What is the center and radius of circle S?

OpenStudy (anonymous):

@Calcmathlete

OpenStudy (anonymous):

General equation of circle is given as: \[(x-h)^2 + (y - k) ^2 = r^2\] Can you find h and k by comparing it with the equation given??

OpenStudy (anonymous):

The form of a circle is: \[(x - h)^{2} + (y - k)^2 = r^2\]THe center is (-h, -k) and the radius is r. Can you see if you can figure it out?

OpenStudy (anonymous):

not too sure how to use that

OpenStudy (anonymous):

Alright. \[(x - 16)^2 + (y - 9)^2 = 4\]For the center, set each equal to zero. x - 16 = 0 y - 9 = 0 Can you solve for x and y individually?

OpenStudy (anonymous):

16 and 9

OpenStudy (anonymous):

You have to just compare the two equations.. What is in place of h and k in your equation??

OpenStudy (anonymous):

Yes you are right it is 16 and 9.. So, (16,9) will be the center..

OpenStudy (anonymous):

and the radius is 2?

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

Yes very correct..

OpenStudy (anonymous):

thanks!!!

OpenStudy (anonymous):

Welcome dear..